OpenStudy (anonymous):

tours are starting, the first tour is 15min the second is 20 when will they both be back at the same time?

OpenStudy (anonymous):

try using the calculator

OpenStudy (anonymous):

Hello, I notice this question was not answered. You can solve this by using LCM least Common Multiples. I am going to find the LCM of the two numbers by using prime factorization. 15 = 3 * 5 20 = 2 * 2 * 5 The LCM of the two numbers is 2 * 2 * 3 * 5 Multiply 2 * 2 * 3 * 5 and you will get 60. It will take 60 minutes( 1 hour) before the tours are back at the same time. If you need help with LCM or anything else, post back to this question.